A resistor initially at 300 K has a heat capacity of 10 J K⁻¹ . It is thermally insulated from its surrounding and is connected to a battery from which it takes 1200 J through an electric current. Calculate the change in entropy of (i) the resistor, 

(ii) the surroundings.


1
Expert's answer
2021-04-13T10:54:30-0400

Answer :-


Given:-


"\boxed{T_1=300K}"


"\boxed{C=10JK^{-1}}"


"\boxed{ \Delta Q=1200J}"



We know that


"\boxed{\Delta S=\dfrac{\Delta Q}{T}}"


We get

"\bigstar"

"C=\dfrac{\Delta Q}{\Delta T}\implies \Delta T=\dfrac{\Delta Q}{C}"


"\Delta T=\dfrac{1200J}{10JK^{-1}}=120K"


"T_2=T_1+\Delta T"



"T_2=300K+120K=420K\n\n\u200b\t\n =300K+120K=420K"


"\Delta S=\dfrac{1200J}{420K}=2.86JK^{-1}"



Answer:"\boxed{
\Delta S=2.86JK^{-1}}" 


. entropy of universe remains constant


"\boxed{\Delta S _{resistor} =- \Delta S{surroundings}}"